A plate made from a rectangular sheet of metal x metres by y metres has a corner "rounded" to a quadrant of a circle with radius r.
Show that r=2 sqrt xy-A/4-pi, where A is the area of the "finished" plate, given that r>0.

Answer by mananth(16949)

.
2r(x-2r) +2r(y-2r)+(x-2r)(y-2r) +pi * r^2
2rx-4r^2+2ry-4r^2+xy-2rx-2ry+4r^2 +pir^2=A
-4r^2+xy +pir^2=A
xy-A =4r^2-pir^2
xy-A = r^2(4-pi)